# Bidirectional Gated Mamba for Sequential Recommendation

### v1

@@ 1. Introduction
Recently, Transformer-based models have emerged as the leading approach in sequential recommendation due to their outstanding performance (de Souza Pereira Moreira et al., 2021). By leveraging the powerful self-attention mechanism (Vaswani et al., 2017; Keles et al., 2023), these models have demonstrated a remarkable ability to deliver accurate predictions. However, despite their impressive performance, current transformer-based models are proven inefficient since the amount of computation grows quadratically as the length of the input sequence increases (Keles et al., 2023). Other approaches, such as RNN-based models e.g., GRU4Rec (Jannach and Ludewig, 2017) and MLP-based models e.g., MLP4Rec (Li et al., 2022), are proven to be efficient due to their linear complexity. Nevertheless, they have struggled with handling long and complex patterns (Yoon and Jang, 2023). All these methods above seem to have suffered from a significant trade-off between effectiveness and efficiency. Consequently, a specially designed State Space Model (SSM) called Mamba (Gu and Dao, 2023) has been proposed. By employing simple input-dependent selection on the original SSM (Liu et al., 2024; Hamilton, 1994), it has demonstrated remarkable efficiency and effectiveness.
@@
  • Context Modeling: While previous research has demonstrated Mamba’s reliability in capturing sequential information (Gu and Dao, 2023; Yang et al., 2024), its unidirectional architecture imposes significant limitations in SRS applications. By only capturing users’ past behaviors, Mamba cannot leverage future contextual information, potentially leading to an incomplete understanding of user preferences  (Liu et al., 2024; Sun et al., 2019). For instance, if a user consistently purchases household items but begins to show interest in sports equipment, a model that does not consider future contexts may struggle to recognize this shift, resulting in sub-optimal next-item predictions (Jiang et al., 2024; Kweon et al., 2021).
  • Short Sequence Modeling: This challenge is primarily driven by the long-tail user problem, a common issue in sequential recommendation. Long-tail users refer to such users who interact with only a few items but typically receive lower-quality recommendations compared to the normal ones (Kim et al., 2019a, b). Furthermore, the instability in state estimation caused by limited data in short sequences (Gu and Dao, 2023; Smith et al., 2022; Yu et al., 2024) exacerbates this problem when Mamba is directly applied to SRS, highlighting the need for effective short sequence modeling. For illustration, we compare two leading baselines, Mamba4Rec (Liu et al., 2024) and SASRec (Kang and McAuley, 2018), against our proposed framework on the Beauty dataset. As shown in Figure 1, the histogram depicts the number of users in each group, while the line represents recommendation performance in terms of Hit@10. SASRec outperforms Mamba4Rec in the first three groups, demonstrating Mamba4Rec’s exacerbation of the long-tail user problem.
To address these challenges and better leverage Mamba’s strengths, we propose an innovative framework called SelectIve Gated MAmba for Sequential Recommendation (SIGMA). Our approach introduces the Partially Flipped Mamba (PF-Mamba), a specialized bidirectional structure that captures contextual information (Liu et al., 2024; Jiang et al., 2024). We then introduce an input-dependent Dense Selective Gate (DS Gate) to allocate the weights of the two directions and further filter the information. Additionally, we develop a Feature Extract GRU (FE-GRU) to better model short-term patterns in interaction sequences (Hidasi et al., 2015), offering a possible solution to the long-tail user problem. Our contributions are summarized as follows::
@@ 2.2. Embedding Layer
For existing SRSs, It is necessary to map the sequential information in user-item interaction to a high-dimensional space (Liu et al., 2023) to effectively capture the temporal dependencies. In our framework, we choose a commonly used method for constructing the item embedding. Here, we denote the user set as $\mathcal{U}=\left\{{u}_{1},{u}_{2},\cdots,{u}_{\left|\mathcal{U}\right|}\right\}$ and the item set as $\mathcal{V}=\left\{{v}_{1},{v}_{2},\cdots,{v}_{\left|\mathcal{V}\right|}\right\}$. So for a chronologically ordered interaction sequence, it can be expressed as $\boldsymbol{S}_{u}=\left[{s}_{1},{s}_{2},\cdots,{s}_{{n}_{u}}\right]$, where ${n}_{u}$ represents the length of the sequence for user $u\in\mathcal{U}$. For simplicity, we omit the mark $\left(u\right)$ in the following sections. Regarding this interaction sequence as the input tensor, we denote $D$ as the embedding dimension and use a learnable item embedding matrix $\boldsymbol{E}\in\mathbb{{R}^{\mathit{\left|\mathcal{V}\right|\times D}}}$ to adaptively projected ${s}_{i}$ into the representation $\boldsymbol{h}_{i}$. The whole interaction sequence can be output as:
@@ 2.4. PFFN Network
To capture the complex features, we further leverage a position-wise feed-forward network (PFFN Net) (Liu et al., 2024; Kang and McAuley, 2018):
@@ 3.1.3. Implementation Details
In this section, we provide a detailed description of our framework’s implementation. The code can be found in Supplementary Material. For GPU selection, all experiments are conducted on a single NVIDIA L4 GPU. The Adam optimizer (Kingma and Ba, 2014) is used with a learning rate of 0.001. For a fair comparison, the embedding dimension for all tested models is set to 64. Other implementation details are the same as original papers (Liu et al., 2024; Wang et al., 2024; Kang and McAuley, 2018), which are presented in Appendix D.
@@ 3.1.4. Baselines
To demonstrate the effectiveness and efficiency of our proposed framework, we compare SIGMA with state-of-the-art transformer-based models (BERT4Rec (Sun et al., 2019), SASRec (Kang and McAuley, 2018), LinRec (Liu et al., 2023), FEARec (Du et al., 2023)), RNN-based models (GRU4Rec (Jannach and Ludewig, 2017)), and SSM-based models (Mamba4Rec (Liu et al., 2024), denoted as Mamba, ECHOMamba4Rec (Wang et al., 2024)), denoted as ECHO. Additional details are provided in Appendix E.
@@ 4.1. Sequential Recommendation
Advancements in deep learning have transformed recommendation systems, making them more personalized and accurate in next-item prediction. Early sequential recommendation frameworks have adopted CNNs and RNNs to capture users’ preferences but faced issues like catastrophic forgetting when dealing with long-term dependencies (de Souza Pereira Moreira et al., 2021; Kim et al., 2019a). Then, the transformer-based models have emerged as powerful methods with their self-attention mechanism, significantly improving performance by selectively capturing the complex user-item interactions (Kang and McAuley, 2018). However, they have suffered from inefficiency due to the quadratic computational complexity (Keles et al., 2023). Therefore, to address the trade-off between effectiveness and efficiency, we propose SIGMA, a novel framework that achieves remarkable performance.

### v2

@@ 1. Introduction
Recently, Transformer-based models have emerged as the leading approach in sequential recommendation due to their outstanding performance (de Souza Pereira Moreira et al., 2021). By leveraging the powerful self-attention mechanism (Vaswani et al., 2017; Keles et al., 2023), these models have demonstrated a remarkable ability to deliver accurate predictions. However, despite their impressive performance, current transformer-based models are proven inefficient since the amount of computation grows quadratically as the length of the input sequence increases (Keles et al., 2023). Other approaches, such as RNN-based models e.g., GRU4Rec (Jannach and Ludewig, 2017) and MLP-based models e.g., MLP4Rec (Li et al., 2022a), are proven to be efficient due to their linear complexity. Nevertheless, they have struggled with handling long and complex patterns (Yoon and Jang, 2023). All these methods above seem to have suffered from a significant trade-off between effectiveness and efficiency. Consequently, a specially designed State Space Model (SSM) called Mamba (Gu and Dao, 2023) has been proposed. By employing simple input-dependent selection on the original SSM (Liu et al., 2024a; Hamilton, 1994), it has demonstrated remarkable efficiency and effectiveness.
@@
  • Context Modeling: While previous research has demonstrated Mamba’s reliability in capturing sequential information (Gu and Dao, 2023; Yang et al., 2024), its unidirectional architecture imposes significant limitations in SRS applications. By only capturing users’ past behaviors, Mamba cannot leverage future contextual information, potentially leading to an incomplete understanding of user preferences  (Liu et al., 2024a; Sun et al., 2019). For instance, if a user consistently purchases household items but begins to show interest in sports equipment, a model that does not consider future contexts may struggle to recognize this shift, resulting in sub-optimal next-item predictions (Jiang et al., 2024; Kweon et al., 2021).
  • Short Sequence Modeling: This challenge is primarily driven by the long-tail user problem, a common issue in sequential recommendation. Long-tail users refer to such users who interact with only a few items but typically receive lower-quality recommendations compared to the normal ones (Kim et al., 2019a, b). Furthermore, the instability in state estimation caused by limited data in short sequences (Gu and Dao, 2023; Smith et al., 2022; Yu et al., 2024) exacerbates this problem when Mamba is directly applied to SRS, highlighting the need for effective short sequence modeling. For illustration, we compare two leading baselines, Mamba4Rec (Liu et al., 2024a) and SASRec (Kang and McAuley, 2018), against our proposed framework on the Beauty dataset. As shown in Figure 1, the histogram depicts the number of users in each group, while the line represents recommendation performance in terms of Hit@10. SASRec outperforms Mamba4Rec in the first three groups, demonstrating Mamba4Rec’s exacerbation of the long-tail user problem.
To address these challenges and better leverage Mamba’s strengths, we propose an innovative framework called SelectIve Gated MAmba for Sequential Recommendation (SIGMA). Our approach introduces the Partially Flipped Mamba (PF-Mamba), a specialized bidirectional structure that captures contextual information (Liu et al., 2024a; Jiang et al., 2024). We then introduce an input-dependent Dense Selective Gate (DS Gate) to allocate the weights of the two directions and further filter the information. Additionally, we develop a Feature Extract GRU (FE-GRU) to better model short-term patterns in interaction sequences (Hidasi et al., 2015), offering a possible solution to the long-tail user problem. Our contributions are summarized as follows::
@@ 2.2. Embedding Layer
For existing SRSs, It is necessary to map the sequential information in user-item interaction to a high-dimensional space (Liu et al., 2023a) to effectively capture the temporal dependencies. In our framework, we choose a commonly used method for constructing the item embedding. Here, we denote the user set as $\mathcal{U}=\left\{{u}_{1},{u}_{2},\cdots,{u}_{\left|\mathcal{U}\right|}\right\}$ and the item set as $\mathcal{V}=\left\{{v}_{1},{v}_{2},\cdots,{v}_{\left|\mathcal{V}\right|}\right\}$. So for a chronologically ordered interaction sequence, it can be expressed as $\boldsymbol{S}_{u}=\left[{s}_{1},{s}_{2},\cdots,{s}_{{n}_{u}}\right]$, where ${n}_{u}$ represents the length of the sequence for user $u\in\mathcal{U}$. For simplicity, we omit the mark $\left(u\right)$ in the following sections. Regarding this interaction sequence as the input tensor, we denote $D$ as the embedding dimension and use a learnable item embedding matrix $\boldsymbol{E}\in\mathbb{{R}^{\mathit{\left|\mathcal{V}\right|\times D}}}$ to adaptively projected ${s}_{i}$ into the representation $\boldsymbol{h}_{i}$. The whole interaction sequence can be output as:
@@ 2.4. PFFN Network
To capture the complex features, we further leverage a position-wise feed-forward network (PFFN Net) (Liu et al., 2024a; Kang and McAuley, 2018):
@@ 3.1.3. Implementation Details
In this section, we provide a detailed description of our framework’s implementation. The code can be found in Supplementary Material. For GPU selection, all experiments are conducted on a single NVIDIA L4 GPU. The Adam optimizer (Kingma and Ba, 2014) is used with a learning rate of 0.001. For a fair comparison, the embedding dimension for all tested models is set to 64. Other implementation details are the same as original papers (Liu et al., 2024a; Wang et al., 2024; Kang and McAuley, 2018), which are presented in Appendix D.
@@ 3.1.4. Baselines
To demonstrate the effectiveness and efficiency of our proposed framework, we compare SIGMA with state-of-the-art transformer-based models (BERT4Rec (Sun et al., 2019), SASRec (Kang and McAuley, 2018), LinRec (Liu et al., 2023a), FEARec (Du et al., 2023)), RNN-based models (GRU4Rec (Jannach and Ludewig, 2017)), and SSM-based models (Mamba4Rec (Liu et al., 2024a), denoted as Mamba, ECHOMamba4Rec (Wang et al., 2024)), denoted as ECHO. Additional details are provided in Appendix E.
@@ 4.1. Sequential Recommendation
Advancements in deep learning have transformed recommendation systems, making them more personalized and accurate in next-item prediction (Liu et al., 2023c, b, 2024b; Li et al., 2022b; Gao et al., 2024; Li et al., 2023a, b; Chen et al., 2023; Cui et al., 2024). Early sequential recommendation frameworks have adopted CNNs and RNNs to capture users’ preferences but faced issues like catastrophic forgetting when dealing with long-term dependencies (de Souza Pereira Moreira et al., 2021; Kim et al., 2019a). Then, the transformer-based models have emerged as powerful methods with their self-attention mechanism, significantly improving performance by selectively capturing the complex user-item interactions (Kang and McAuley, 2018). However, they have suffered from inefficiency due to the quadratic computational complexity (Keles et al., 2023). Therefore, to address the trade-off between effectiveness and efficiency, we propose SIGMA, a novel framework that achieves remarkable performance.